### v1

# Non-stabilizerness in open XXZ spin chains: Universal scaling and dynamics

Abstract
Magic, or non-stabilizerness, is a crucial quantum resource, yet its dynamics in open quantum systems remain largely unexplored. We investigate magic in the open XXZ spin chain under either boundary gain and loss, or bulk dephasing using the stabilizer Rényi entropy $M_{2}$. To enable scalable simulations of large systems, we develop a novel, highly efficient algorithm for computing $M_{2}$ within the matrix product states formalism while maintaining constant bond dimension—an advancement over existing methods. For boundary driving, we uncover universal scaling laws, $M_{2}(t)\sim t^{1/z}$, linked to the dynamical exponent $z$ for several distinct universality classes. We also disentangle classical and quantum contributions to magic by introducing a mean-field approximation for magic, thus emphasizing the prominent role of quantum critical fluctuations in non-stabilizerness. For bulk dephasing, dissipation can transiently enhance magic before suppressing it, and drive it to a nontrivial steady-state value. These findings position magic as a powerful diagnostic tool for probing universality and dynamics in open quantum systems.
Introduction.
Non-stabilizerness quantifies the deviation of a quantum state from the set of stabilizer states, serving as a key resource for quantum advantage [1, 2, 3, 4] in tasks such as quantum computing and simulation [5, 6, 7, 8]. In closed systems, unitary evolution generates and propagates magic through coherent interference, with chaotic dynamics typically leading to faster magic growth compared to localized phases [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. However, in open quantum systems, the interplay between coherent dynamics and decoherence fundamentally alters the behavior of magic. While dissipation and dephasing generally suppress magic by destroying quantum coherences, driven-dissipative systems can exhibit non-equilibrium steady states (NESS) that preserve or even generate magic, particularly in the presence of interactions [19, 20, 21]. Understanding the dynamics of magic in open systems is therefore crucial for harnessing quantum resources in realistic, noisy environments [22, 23, 24].
In this work, we investigate the fate of magic in open quantum systems, focusing on the XXZ spin chain under boundary driving and bulk dephasing [25, 26, 27, 28, 29, 30]. Using the stabilizer Rényi entropy $M_{2}$ as a measure of magic [31], we explore how decoherence and dissipation reshape the Pauli spectrum and influence the growth and saturation of magic.
First, we develop a highly efficient algorithm for computing the stabilizer Rényi entropy (magic) in open quantum systems, which maintains a constant bond dimension throughout the calculation. This represents a significant improvement over existing methods, as it avoids the bond dimension growth typically associated with tensor network approaches [32, 33, 34].
For boundary gain and loss, we identify universal scaling law for magic in the XXZ chain, showing $M_{2}(t)\sim t^{1/z}$ with $z$ the dynamical critical exponent. It reflects various transport regimes, including ballistic $(z=1)$, diffusive $(z=2)$, and even the Kardar-Parisi-Zhang (KPZ) dynamics [35, 36, 37] at the isotropic point with $z=3/2$. To disentangle classical and quantum contributions to non-stabilizerness, we introduce a mean-field approximation for magic that eliminates quantum correlations while preserving local magnetization contributions. This approach provides a classical baseline for magic, allowing us to quantify the role of quantum correlations, and identify their impact on transport dynamics and magic generation.
In the presence of bulk dephasing when all sites are coupled to environment through coupling $\gamma_{z}$, we find purely dissipative dynamics with no coherent evolution at the spin rotational invariant limit. This occurs when starting from a fully polarized state in any direction. In general, we find that the time evolution is governed by two distinct timescales: the dephasing timescale $\tau_{d}\sim 1/\gamma_{z}$, which drives the exponential suppression of magic, and the coherence timescale $\tau_{c}$. The latter is the inverse of the SU(2) symmetry breaking interaction and controls transient oscillations in magic. Remarkably, dephasing can transiently enhance magic before driving it to zero. However, in magnetization-conserving systems, magic does not fully decay but stabilizes at a non-trivial steady-state value. Furthermore, when the dynamics is restricted to the zero magnetization sector, starting from certain initial states such as the Néel configuration, magic exhibits a robust power-law decay over time, reflecting the underlying gapless nature of the Lindbladian [38, 39].
Stabilizer Rényi entropy in open systems.
Figure 1: (a) Illustration of the MPS structure for the vectorized density matrix, with explicit site labels shown. (b) Depiction of the Pauli vector construction and its corresponding MPS structure. The red dots represent the Pauli tensors. (c) After contracting the physical indices, the Pauli vector is expressed as an MPS with the same bond dimension $\chi$ as the original vectorized density matrix.
In open quantum systems, magic can be extended to mixed states through the stabilizer Rényi entropy, defined using the density matrix [31]. For a general state $\rho$, the stabilizer Rényi entropy $M_{2}$ is given by
$M_{2}=-\log_{2}\frac{\sum_{P}|c_{P}|^{4}}{\sum_{P}|c_{P}|^{2}},$ (1)
where $P$ represents elements of the $N$-qubit Pauli group [40, 41, 42], and the Pauli coefficients $c_{P}$ are $c_{P}=\mathrm{Tr}(\rho P)$. In open systems, the Pauli spectrum exhibits both delocalization due to coherent dynamics and decay from non-unitary processes. Even for closed systems, calculating the full Pauli spectrum is computationally expensive for large $N$, due to the exponential growth of the Pauli group. Efficient methods like Monte Carlo sampling [43] and tensor networks [44, 32, 45] allow for practical evaluation of $M_{2}$ in closed systems.
We introduce a new approach to compute magic in open systems, using a vectorized density matrix structure while maintaining a constant bond dimension. This method, extending the framework of Ref. [45], avoids the bond dimension increase required by closed-system algorithms, offering a scalable and computationally efficient tool for quantifying magic in open quantum systems.
MPS in the vectorized basis.
Figure 2: (a,b,c) Time evolution of magic for different anisotropy parameters $\Delta$ and for various system sizes. The insets in each panel present the light cone formation in magnetization for the respective $\Delta$ as a function of position $x=2j/L-1$, with $j$ numbering the sites. The initial density matrix corresponds to the infinite-temperature state, $|\rho_{0}\rangle\!\rangle=|\mathbb{1}_{\infty}\rangle\!\rangle$. For $\Delta=0$, the magic and magnetization light cone exhibit a linear growth, characteristic for the ballistic transport regime. At $\Delta=1$, the growth follows a $t^{2/3}$ scaling, indicative of the KPZ universality class, while for $\Delta=2$, it becomes diffusive with a dynamical exponent $z=2$. The legend and colorbar are shared for all panels.
One of the main results of this work is the development of a highly efficient algorithm for computing stabilizer Rényi entropy magic in quantum systems, which maintains a constant bond dimension throughout the calculation, ensuring scalability and computational efficiency even for large systems. We represent the vectorized density matrix $|\rho\rangle\!\rangle$ in matrix product state (MPS) form [46, 47, 48]
$|\rho\rangle\!\rangle=\sum_{s_{i},s^{\prime}_{i}}A_{1}^{s_{1},s^{\prime}_{1}}A_{2}^{s_{2},s^{\prime}_{2}}\cdots A_{L}^{s_{L},s^{\prime}_{L}}|s_{1}s^{\prime}_{1}\dots s_{L}s^{\prime}_{L}\rangle,$ (2)
where we explicitly track the pairs of physical indices $(s_{i},s^{\prime}_{i})$ at each site. The tensors $A_{j}^{s_{j},s^{\prime}_{j}}$ are of dimension $\chi\times\chi$ for $1<j<L$, while the boundary tensors $A_{1}^{s_{1},s^{\prime}_{1}}$ and $A_{L}^{s_{L},s^{\prime}_{L}}$ are a $1\times\chi$ row vector and a $\chi\times 1$ column vector, respectively. This structure is illustrated in Fig. 1. The normalization condition $\mathrm{Tr}(\rho)=1$ translates to $\langle\!\langle\mathbb{1}_{\infty}|\rho\rangle\!\rangle=1$, where $|\mathbb{1}_{\infty}\rangle\!\rangle$ represents the MPS form of the infinite-temperature density matrix. To compute the stabilizer Rényi entropy, we define the Pauli vector $|P\rho\rangle$ by applying a tensor product of Pauli operators:
$P_{\bm{\alpha}}=P_{\alpha_{1}}\otimes P_{\alpha_{2}}\otimes\dots\otimes P_{\alpha_{L}},$ (3)
and the corresponding vectorized form is
$|P\rho\rangle=\sum_{\alpha_{i}}B_{1}^{\alpha_{1}}B_{2}^{\alpha_{2}}\cdots B_{L}^{\alpha_{L}}|\alpha_{1}\dots\alpha_{L}\rangle.$ (4)
The tensors $B_{j}^{\alpha_{j}}$ are defined as
$B_{j}^{\alpha_{j}}=\sum_{s_{j},s^{\prime}_{j}}\frac{\langle s_{j}|P_{\alpha_{j}}|s^{\prime}_{j}\rangle}{\sqrt{2}}A_{j}^{s_{j},s^{\prime}_{j}}.$ (5)
Crucially, the bond dimensions of $|P\rho\rangle$ are identical to those of $|\rho\rangle\!\rangle$, ensuring that the computational cost remains manageable. The Pauli vector is normalized such that $\langle P\rho|P\rho\rangle=\sum_{\bm{\alpha}}\mathrm{Tr}(P_{\bm{\alpha}}\rho)^{2}$. Next, we introduce a matrix product operator (MPO) for efficient computation:
$W=\sum_{\alpha_{i},\alpha^{\prime}_{i}}C_{1}^{\alpha_{1},\alpha_{1}^{\prime}}C_{2}^{\alpha_{2},\alpha^{\prime}_{2}}\cdots C_{L}^{\alpha_{L},\alpha^{\prime}_{L}}|\alpha_{1}\dots\alpha_{L}\rangle\langle\alpha^{\prime}_{1}\dots\alpha^{\prime}_{L}|.$ (6)
The MPO tensors are given by $C_{j}^{\alpha_{j},\alpha^{\prime}_{j}}=B_{j}^{\alpha_{j}}\delta_{\alpha_{j},\alpha^{\prime}_{j}}$, and this construction allows for the computation of the stabilizer Rényi entropy (magic) as
$M_{2}=-\log_{2}\frac{\langle P\rho|W^{2}|P\rho\rangle}{\langle P\rho|P\rho\rangle}.$ (7)
In practice, we apply $W$ to $|P\rho\rangle$ and compute its inner product with itself to obtain the numerator of Eq. (7). The constant bond dimension throughout this process is a central feature of the method, making it a powerful and scalable tool for analyzing quantum systems, particularly in the context of stabilizer entropy.
Boundary driven XXZ spin chain.
The open XXZ spin chain is a paradigmatic model for studying transport and nonequilibrium steady states in open quantum systems [49, 26, 21]. Its Hamiltonian is given by
$H=\frac{J}{4}\sum_{j=1}^{L-1}\left(X_{j}X_{j+1}+Y_{j}Y_{j+1}+\Delta Z_{j}Z_{j+1}\right),$ (8)
where $X_{j},Y_{j}$, and $Z_{j}$ are Pauli operators at site $j$, and $\Delta$ is the anisotropy parameter. We investigate the generation and depletion of magic in the XXZ chain coupled to external reservoirs. Specifically, we focus on boundary-driven dynamics, where interactions with reservoirs at the chain’s edges drive the system into a nonequilibrium steady state (NESS) [25, 49] as displayed in Fig. 3(a). The dissipation at the boundaries is described by the Lindblad operators
$F_{1,1}$ $=\sqrt{\gamma}\sqrt{1-\mu}\,S_{1}^{-},$ $F_{1,2}$ $=\sqrt{\gamma}\sqrt{1+\mu}\,S_{1}^{+},$
$F_{L,1}$ $=\sqrt{\gamma}\sqrt{1+\mu}\,S_{L}^{-},$ $F_{L,2}$ $=\sqrt{\gamma}\sqrt{1-\mu}\,S_{L}^{+},$ (9)
which act at the first and last sites of the chain. The system is initialized in an infinite-temperature density matrix, $|\rho_{0}\rangle\!\rangle=|\mathbb{1}_{\infty}\rangle\!\rangle$, a stabilizer state with zero magic. At $t=0$, the couplings to the reservoirs are switched on, and we track the time evolution of the magic measure $M_{2}(t)$, along with the local magnetization profile $\langle S_{z}(x,t)\rangle$. In our numerical simulations, the driving parameter is kept small, $0.01\leq\mu\leq 0.05$.
Following the quench, the system undergoes a dynamical evolution, eventually reaching a nonequilibrium steady state (NESS) characterized by a constant magnetization current [49]. The emergence of two expanding light cones, originating from the system’s boundaries, is clearly visible in the magnetization profiles of Fig. 2(a,b,c), where their spatial spread follows the dynamical scaling $x\sim t^{1/z}$. The exponent $z$ dictates the transport behavior, distinguishing different universality classes. A crucial aspect of this evolution is the initial state choice. The system is initialized in the infinite-temperature density matrix, $|\rho_{0}\rangle\!\rangle=|\mathbb{1}_{\infty}\rangle\!\rangle$, which ensures an uncorrelated and homogeneous starting point. This setup allows the observed transport properties and the associated dynamical scaling to emerge purely from the interplay between unitary evolution and boundary driving, rather than from any pre-existing correlations. In particular, it guarantees that the KPZ universality class at $\Delta=1$ arises intrinsically from the system’s dynamics, reinforcing the robustness of the observed scaling behavior.
For $\Delta<1$, the system exhibits ballistic transport [50], with a linear light cone ($z=1$). At the critical point $\Delta=1$, the dynamics enters the Kardar-Parisi-Zhang (KPZ) universality class [37, 51, 52], where transport becomes superdiffusive with a characteristic light cone scaling $z=3/2$. This anomalous behavior reflects the emergence of strong correlations and fluctuations. In contrast, for $\Delta>1$, transport is diffusive ($z=2$), resulting in a broader and slower-growing light cone.
Remarkably, the time evolution of magic follows the same universal scaling as the magnetization profile, with $M_{2}(t)\sim t^{1/z}$. Rescaling by $\mu^{2}$ and by the system size $L$ reveals a universal collapse of the magic dynamics across different transport regimes, underscoring its role as a robust measure of the system’s universality class. In the KPZ regime ($\Delta=1$), magic exhibits a characteristic $t^{2/3}$ growth, in precise agreement with the expected scaling. This highlights magic as a sensitive probe of quantum transport, capturing the emergence of quantum correlations that define the KPZ universality class, beyond what is accessible through conventional observables like magnetization currents. Similar to the magnetization profile, magic does not develop uniformly but instead emerges within the two light cones propagating from the system’s edges. These expanding regions serve as the primary sources for magic generation before it eventually saturates in the NESS. As shown in Fig. 2 (insets), the transition to the steady state coincides with the collision of the two light cones at the center, marking the completion of the transport-driven magic buildup.
One of the key findings of this work is the emergence of a universal scaling law for magic, as demonstrated in Fig. 2 (d,e,f). Our results establish that the magic follows the scaling form
$M_{2}(L,t)\sim L\,f\left(\frac{t}{L^{z}}\right),$ (10)
where the scaling function $f(x)$ governs the dynamical evolution across different regimes. In the early-time regime ($t\ll L^{z}$), magic exhibits a universal growth $M_{2}(t)\sim t^{1/z}$, directly reflecting the system’s transport properties, $f(x\ll 1)\sim x^{1/z}$. At late times ($t\gg L^{z}$), magic saturates as $M_{2}\sim L$, signaling the approach to the nonequilibrium steady state, which gives $f(x\gg 1)\sim const$ for the scaling function. This universal scaling structure provides a unifying framework for understanding its growth across different dynamical universality classes.
Figure 3: (a) Sketch of the boundary driven XXZ spin chain in the infinite temperature limit. (b) The steady-state values of the total magic, $M_{2}(t_{\infty})$, and the mean-field magic, $M_{2}^{\rm{MF}}(t_{\infty})$, as functions of $\Delta$. The hashed region represents the contribution to magic arising from quantum correlations. (c) Steady-state magnetization profile $S_{z}(t_{\infty})$ for three different values of $\Delta$. In both panels, the system size is set to $L=32$.
Figure 4: (a) Sketch of the bulk dephasing modeled with sites coupled to an external reservoir via the jump operator $S^{z}$. (b, c) Time evolution under dephasing affecting all qubits at the isotropic point $\Delta=1$. (b) Starting from an initial density matrix with zero magic, $\rho(0)=(|+\rangle\langle+|)^{\otimes L}$, the magic initially increases before undergoing an exponential decay to zero at late times. The inset displays the magic density in the NESS state $M_{2}(t_{\infty},\theta)$ as a function of the initial polarization angle $\theta$. (c) When the initial density matrix is $\rho(0)=(|T\rangle\langle T|)^{\otimes L}$, the system possesses a high degree of magic. Due to conservation of the total $S_{z}$ spin component there is a non-zero magic in the NESS $M_{2}(t_{\infty})=\log_{2}(6/5)$, subtracted in $(b)$ from $M_{2}(t)$. The dashed line denotes the analytical result from Eq. (11).
Mean-field approximation for magic.
To quantify the quantum nature of the system, we introduce a mean-field approximation that isolates classical contributions to magic. The mean-field density matrix, $\rho_{\rm MF}(t)$, is a product state with bond dimension one, constructed to reproduce the local magnetization while discarding all correlations: $\rho_{\rm MF}\approx\bigotimes_{i=1}^{L}\rho_{i}$, where each local density matrix $\rho_{i}$ is given by $\rho_{i}=\frac{1}{2}\left(\mathbb{1}_{2}+\sum_{\alpha}m_{i}^{\alpha}\sigma_{i}^{\alpha}\right)$. Here, $\mathbb{1}_{2}$ is the identity matrix, $\sigma_{i}^{\alpha}$ ($\alpha=x,y,z$) are the Pauli matrices at site $i$, and $m_{i}^{\alpha}=\langle\sigma_{i}^{\alpha}\rangle$ are the local magnetization components. This formulation allows us to compute the mean-field magic $M_{2}^{\rm{MF}}(t)$ via Eq. (1) and compare it directly with the full magic $M_{2}(t)$, thereby quantifying the contribution of quantum correlations.
Figure 3 (b) highlights the fundamental role of transport in determining the steady-state magic. In the ballistic regime ($\Delta\ll 1$), quantum correlations are dominant, maximizing the gap between $M_{2}(t_{\infty})$ and $M_{2}^{\rm{MF}}(t_{\infty})$. Since the magnetization profile remains nearly unchanged for $\Delta\leq 0.6$ [see Fig. 3 (c)], the mean-field magic stays almost constant. The small amplitude of $M_{2}^{\rm{MF}}(t_{\infty})$ reflects the local magnetization, which almost vanishes in the bulk, leaving boundary sites as the primary contributors. Near $\Delta\approx 1$, both measures of magic rise sharply, signaling enhanced entanglement and a fundamental shift in the magnetization profile. In the diffusive regime, the gap between full and mean-field magic narrows, reflecting the suppression of quantum correlations.
The explicit decomposition of magic into classical and quantum components is a key achievement of this work, as it provides a direct measure of non-stabilizerness due to quantum correlations. Notably, despite being entirely classical, $M_{2}^{\rm{MF}}(t)$ exhibits the same scaling behavior as the full magic [39], confirming that universal transport signatures persist even in the absence of quantum correlations.
Quantum Dephasing.
We investigate bulk dephasing using local Lindblad operators $F_{j}=\sqrt{\gamma_{z}}S_{j}^{z}$, which act independently at each site [see Fig. 4(a)], suppressing quantum coherences in the $S^{z}$ basis and leading to information loss. To understand its effects, we first analyze the single-qubit case. The density matrix evolves as $\rho_{1}(\theta,\varphi,t)=\frac{1}{2}\left[1+Z\cos\theta+\sin\theta(X\cos\varphi+Y\sin\varphi)e^{-\gamma_{z}t/2}\right]$, where dephasing exponentially suppresses $X$ and $Y$ while preserving $Z$. The corresponding magic follows
$M_{2}(t)=\log_{2}\frac{1+\cos^{2}\theta+\sin^{2}\theta e^{-\gamma_{z}t}}{1+\cos^{4}\theta+\sin^{4}\theta(\sin^{4}\varphi+\cos^{4}\varphi)e^{-2\gamma_{z}t}}.$ (11)
This highlights how dephasing erases quantum coherences while retaining classical information along $Z$. A similar analysis extends to two interacting qubits, leading to a general expression for magic (see Ref. [39] for details).
Heisenberg chain.
In the Heisenberg limit ($\Delta=1$), full $SU(2)$ spin symmetry ensures that any fully polarized product state remains an eigenstate, leading to purely dephasing dynamics with no unitary evolution. As a result, each qubit evolves independently, and the magic density follows the single-qubit result. However, away from this special point, two competing timescales emerge: the dephasing time $\tau_{d}\sim 1/\gamma_{z}$, governing the exponential suppression of magic, and the coherence time $\tau_{c}\sim 1/J|\Delta-1|$, associated with coherent oscillations in magic (see Ref. [39] for details). At $\Delta=1$, the absence of coherence-driven dynamics eliminates $\tau_{c}$, leading to a universal decay dictated solely by $\tau_{d}$. Figure 4 shows the time evolution for two initial states. For $\rho_{0}=(|+\rangle\langle+|)^{\otimes L}$ [Fig. 4(a)], magic starts at zero, peaks at $t\approx 1/\gamma_{z}$, and then decays exponentially. This transient magic enhancement reflects the interplay between decoherence timescales. When starting from a high-magic state $\rho_{0}=(|T\rangle\langle T|)^{\otimes L}$ [Fig. 4(b)], the conservation of $\langle S_{z}\rangle$ prevents full decay, stabilizing a nontrivial steady-state magic,
$\lim_{t\to\infty}\frac{M_{2}(t)}{L}=\log_{2}\frac{1+\cos^{2}\theta}{1+\cos^{4}\theta}.$ (12)
When rescaled by system size, magic dynamics collapse onto a universal curve, matching the single-qubit behavior. The inset of Fig. 4(a) shows a polar plot of the steady-state magic.
Conclusions.
In this work, we have uncovered the previously unexplored dynamics of magic in open quantum systems. By developing a novel and efficient algorithm for computing the stabilizer Rényi entropy $M_{2}$ in open systems, we achieve scalable simulations of large systems under non-unitary dynamics.
For boundary driving, we reveal a universal scaling law for magic, $M_{2}(t)\sim t^{1/z}$, directly tied to the dynamical exponent $z$ of the underlying transport regime, and demonstrate that magic serves as a sensitive probe of quantum transport, including the Kardar-Parisi-Zhang universality class.
We introduce a magic mean-field approach, which isolates the classical contribution to magic through local expectation values. The difference between full and magic mean-field serves as a direct probe of non-stabilizerness generated by the many-body correlations built into the system.
For bulk dephasing, we demonstrate that dissipation can transiently enhance magic before ultimately driving it toward its steady state value. Magic exhibits an initial growth, peaks around a characteristic dephasing timescale $t\approx 1/\gamma_{z}$, and then decays exponentially. In systems with conserved magnetization, however, magic is not fully erased and stabilizes at a non-trivial steady-state value. In the Heisenberg limit, we derive an analytical expression capturing this time dependence. Away from the SU(2) point, the dynamics are governed by two key timescales: the dephasing time $\tau_{d}\sim 1/\gamma_{z}$, which controls the exponential decay, and the coherence time $\tau_{c}\sim 1/J|\Delta-1|$, which accounts for transient oscillations away from the $SU(2)$ point. Intriguingly, when the system is initialized in e.g. the Néel or $(|+\rangle\langle+|)^{\otimes L}$ configurations within the zero-magnetization sector—magic decays as a power law in time, revealing signatures of the gapless nature of the Lindbladian.
These results establish magic as a powerful and versatile measure for characterizing quantum dynamics in open systems and highlight the complex interplay between unitary and non-unitary dynamics as well as symmetry in open quantum systems.
Acknowledgements.
This work received financial support from CNCS/CCCDI-UEFISCDI, under projects number PN-IV-P1-PCE-2023-0159, PN-IV-P1PCE-2023-0987, and PN-IV-P8-8.3-PM-RO-FR-2024-0059, and by the “Nucleu” Program within the PNCDI 2022-2027, carried out with the support of MEC, project no. 27N/03.01.2023, component project code PN 23 24 01 04.. This research was also supported by the National Research, Development and Innovation Office - NKFIH within the Quantum Technology National Excellence Program (Project No. 2017-1.2.1-NKP-2017-00001), K134437, K142179 by the BME-Nanotechnology FIKP grant (BME FIKP-NAT), the QuantERA ‘QuSiED’ grant No. 10101773. We acknowledge the Digital Government Development and Project Management Ltd. for awarding us access to the Komondor HPC facility based in Hungary.